OpenStudy (anonymous):

1) solve by the elimination method. 4x-2y= 38 4y= 23 - 3x 2) 5x + y= -3 7x-3y= -35

jimthompson5910 (jim_thompson5910):

I'll do the first one to start you off. First, add 3x to both sides in the 2nd equation to get 3x+4y=23. Now double everything in the first equation to get 8x-4y=76. Now add the two equations to get (8x-4y)+(3x+4y)=76+23 which then simplifies to 11x=99. Solve for x to get x = 99/11 = 9. So x = 9. Plug this into either equation and solve for y. I'm going to pick on the equation 4y=23-3x. So 4y=23-3(9) becomes 4y=23-27 and 4y=-4, which means that y = -1. So the solutions are x = 9 and y = -1

OpenStudy (anonymous):

thanks

OpenStudy (gw2011):

1) rewrite 4y=23-3x to get 3x+4y=23 Now you get: 4x-2y=38 3x+4y=23 Multiply the first equation by 2 to get the following: 8x-4y=76 3x+4y=23 Add both equations: 11x=99 x=9 Substitute this value into : 3x+4y=23 3(9)=4y=23 27+4y=23 4y=-4 y=-1 2) Similar situation to (1) above: 5x+y=-3 7x-3y=-35 Multiply the first equation by 3 to get: 15x+3y=-9 7x-3y=-35 Add both equations to get: 22x=-44 x=-2 Substitute this value into: 5x+y=-3 5(-2)+y=-3 -10+y=-3 y=7